A space station shaped like a giant wheel has a radius of 148 m and a moment of inertia of 4.69 ✕ 108 kg · m2. A crew of 150 lives on the rim, and the station is rotating so that the crew experiences an apparent acceleration of 1g. When 100 people move to the center of the station for a union meeting, the angular speed changes. What apparent acceleration is experienced by the managers remaining at the rim? Assume that the average mass of each inhabitant is 65.0 kg.
Answer #1

Initially, I = 4.69 x 108 kg·m² + 150 x 65kg x (148m)² = 6.83 x 108 kg·m²
Later, I = 4.69 x 108 kg·m² + 50 x 65kg x (148m)² = 5.4 x 108 kg·m²

centripetal a = ω²r
9.8 m/s² = ω² x 148m
ω = 0.26 rad/s

Conserve momentum: initial L = final L
6.83 x 108 kg·m²  x 0.26 rad/s = 5.4 x 108 kg·m²  x ω
ω = 0.329 rad/s

Then a = ω²r = (0.329 rad/s)² * 148 m = 16.02 m/s² ≈ 1.63 g